By using log,solve the following equations.a)5^x=7
b)3^x+2=2^x

Answer by Alan3354(69443)

a)5^x=7
x*log(5) = log(7)
x = log(7)/log(5)
-----------------
b)3^x+2=2^x
I think you mean
3^(x+2) = 2^x
(x+2)*log(3) = x*log(2)
x*log(3) + log(9) = x*log(2)
x(log(3) - log(2)) = -log(9)
x = -log(9)/log(1.5)
x =~ -5.419